# Learning to Reason Across Parallel Samples for LLM Reasoning

### v1

@@ 5.1 Comparison with Parallel Scaling Baselines
Figure 2: Compare the performance of SSA RL, PRM, and Majority Vote methods across Qwen 2.5 $\text{LLM}_{\text{ans}}$ model sizes (7B, 14B, 32B) and number of candidate solutions $k=5,10,15,20$.
@@
We present results highlighting the transferability of SSA across datasets, model scales, and model families.
Generalization to harder datasets.
In Table 1, while we have trained the SSA models with problems only from GSM8K and MATH, we observe substantial performance gains on other harder datasets (AMC23 and Olympiad). Such results indicate SSA can generalize to unseen test sets outside of the training datasets.
In addition, we tested SSA on general tasks (specifically ARC-C, MMLU-Pro, and TruthfulQA [5, 44, 24]) without additional training. We observe minor gains compared to majority vote, especially when using both SFT and RL, in ARC-C and MMLU-Pro dataset, but very little for TruthfulQA dataset. Further research is needed to study generalization across different domains. More details can be found in Appendix A.5.
Generalization across scales.
SSA trained on outputs from a smaller answer model (Qwen-2.5-7B) can generalize to outputs from larger answer models (14B, 32B). Figure 2 compare the overall performance of SSA and the other baselines under different numbers of samples. We can see that SSA RL consistently outperform the baselines under different number of candidate solution samples across Qwen models with scales ranging from 7B to 32B. More SSA versions and results are in Table 4 in the Appendix A
Generalization across model families.
In addition to the Qwen 2.5 7B Instruct model, we also tested Llama 3.1 8B Instruct model for inference [9].
@@
Table 2: Results (accuracy %) with $k=5$ candidate from Llama 3.1 8B Instruct $\text{LLM}_{\text{ans}}$ models.
Method Train Data Size Datasets Avg
@@
USC w/ Qwen 7B - 84.15 51.6 6.67 37.5 19.29 39.84
LLama ORM (8B) 273k 89.61 51.8 13.33 30.00 18.4 40.63
Qwen PRM (7B) 3M+ 91.51 56.20 13.33 35.00 20.77 43.36
[…]
SSA RL (1.5B) 140k 88.48 56.60 10.00 27.50 20.47 40.61
SSA RL (3B) 140k 89.08 57.80 10.00 32.50 20.62 42.00
@@
5.3 Comparison with Sequential Test-time Scaling
We have demonstrated the advantageous of SSA over diverse parallel scaling methods in early §5.1. We now compare the performance on SSA and reasoning models that are enabled to spend more test-time compute sequentially with an RL-based approach as in [7]. In particular, we use the results reported by Simple-RL Zoo ([51]) because Zeng et al. [51] also uses Qwen-series models and train them on GSM8K and MATH dataset. In our comparison, we match the test-time computation, specifically, our 3B SSA is trained with Qwen-2.5-7B answers, but we apply it on top of 7B, 14B, 32B $\text{LLM}_{\text{ans}}$ models. We sample 8 candidate solutions from the $\text{LLM}_{\text{ans}}$ model with a maximum token length of 1,024 per answer, roughly matching the maximal generation length, 8,192, of these reasoning models.
@@ 6.2 To think or not to think
A distinct pattern we noted in an RL method is the reduction of the thinking tokens. As shown in Figure 4(b), response length quickly drops during training across all model sizes. Typically, the model often just repeating the provided instruction format for thinking (e.g., ‘<think>reasoning process here</think>’) followed immediately with the final answer.
Examples are shown in Appendix C Figure 14. This contrasts with other RL-based reasoning models that generate longer explanations with more training [7, 51]. It is likely because our SSA is conditioned on multiple candidate solutions, reducing the utility of detailed reasoning.
@@
6.3 Error Analysis
Table 3: Accuracy and error composition for the 3 B SSA ($k{=}5$). Columns are percentages of the whole dataset. $y^{*}\in C$ means ground-truth answer $y^{*}$ included among the 5 candidates $C$.
Acc.$\uparrow$ Correct Wrong
$y^{*}\in C$ copied $y^{*}\notin C$ synthesized $y^{*}\notin C$ copied $y^{*}\notin C$ synthesized $y^{*}\in C$ copied $y^{*}\in C$ synthesized
GSM8K 93.3 87.7 5.5 3.3 1.0 2.3 0.2
MATH 79.2 75.6 3.6 9.4 3.6 6.6 1.2
AMC23 57.5 52.5 5.0 12.5 7.5 12.5 10.0
AIME24 13.3 13.3 0.0 23.3 50.0 3.3 10.0
Olympiad 39.8 34.9 4.9 27.6 24.6 5.6 2.4
7 Discussion and Limitation
Limitations.
Although our approach excels with relatively small inference contexts ($k$), we observed performance degradation at larger contexts, such as $k=20$. This diminishing return might result from several factors, including the limited attention capacity of smaller models or challenges associated with training effectively at larger context lengths. Furthermore, expanding the context length significantly increases VRAM requirements, posing practical constraints on larger $k$ values. It would be difficult to scale $k$ to hundreds of samples like majority vote and PRM due to context length restriction.
Societal Impacts.
Our method’s efficiency in post-training smaller models to improve reasoning performance significantly reduces computational and energy demands compared to directly fine-tuning large-scale models. This decoupling between foundational model development and specialized reasoning enhancement contributes positively to the resource-efficient research.
@@ 8 Conclusion
For future research, promising directions include effectively scaling inference $k$, incorporating different models’ outputs as training data, enhancing the SSA’s synthesis capabilities, extending applications beyond mathematical reasoning, and further reducing SSA model size. Overall, we believe SSA’s hybrid approach offers a promising direction for future research in LLM reasoning.
@@ A.1 SSA Results
We also report the statistical significance of all results. For statistical testing, we employed McNemar’s test, which is appropriate for paired binary outcomes. On our combined dataset analysis (aggregating all 2563 test examples across five benchmarks), our improvements over the baseline are statistically significant (p<0.05) for all configurations.
@@ A.4 Increasing $k$ during training
A.5 Generalization to other tasks
Table 8 reports accuracy on three general tasks benchmarks. On ARC-C and MMLU-Pro the SFT + RL SSA mostly beats the majority-vote baseline [5, 44]. Pure RL SSAs give smaller (sometimes negative) gains, suggesting that an SFT warm-start improves out-of-domain transfer. Pure SFT SSAs exhibit varied performance: strong at 3B scale but significantly weaker at smaller scales (0.5B and 1.5B), suggesting SFT’s generalization capability strongly depends on model capacity for out domain generalization. On TruthfulQA (truthfulness/adversarial) none of the SSAs can consistently outperform majority voting, suggesting that truthfulness might not be beneficial from answer selection. Tackling truthfulness might require different goal than reasoning among answers.
Metric Qwen2.5-7B-Instruct SSA RL SSA SFT + RL SSA SFT
pass@1 Majority Vote (5) 0.5B 1.5B 3B 0.5B 1.5B 3B 0.5B 1.5B 3B
ARC-C 88.65 91.38 88.05 91.3 88.99 91.72 91.38 91.81 69.88 90.02 92.06
A.6 Efficiency Evaluation
Compute Cost (Wall-clock)
Qwen2.5-7B 1 556.12 45.5
Majority vote ($k{=}5$) 5×5$\times$5 ×Qwen2.5-7B 2 780.6 49.67
Qwen PRM 7B ($k{=}5$) 5×5$\times$5 ×Qwen2.5-7B 21.9 2 780.6 + 21.9 = 2 802.5 52.97
SSA RL 3B ($k{=}5$, ours) 5×5$\times$5 ×Qwen2.5-7B 25.7 2 780.6 + 25.7 = 2 806.3 56.13
SimpleRL 7B (seq. RL) 1 pass 3 273.2 58.56
Inference Cost
To measure raw compute (flops), we follow the formula from Kaplan et al. [17]. It uses roughly $C_{forward}\approx 2N$, and the backward pass is about 2 times which is $4N$. So the total is $C\approx 6N$ FLOPs per training token, with $N$ as the model parameter and $C$ as the non embedding training compute. We use $D$ as token length.

### v2

@@ 5.1 Comparison with Parallel Scaling Baselines
Figure 2: Compare the performance of SSA RL, PRM, and Majority Vote methods across Qwen 2.5 $\text{LLM}_{\text{ans}}$ model sizes (7B, 14B, 32B) and number of candidate solutions $k=5,10,15$.
@@
Generalization across scales. SSA trained on outputs from a smaller answer model (Qwen 2.5 7B Instruct) can generalize to outputs from larger answer models (14B, 32B). Figure 2 compare the overall performance of SSA and the other baselines under different numbers of samples. We can see that SSA RL consistently outperform the baselines under different number of candidate solution samples across Qwen models with scales ranging from 7B to 32B. More SSA versions and results are in Table 4 in the Appendix A
Generalization across model families. In addition to the Qwen 2.5 7B Instruct model, we also tested Llama 3.1 8B Instruct model for inference (Grattafiori et al., 2024).
@@
Table 2: Generalization results for different model families (accuracy %) with $k=5$ candidate from Llama 3.1 8B Instruct $\text{LLM}_{\text{ans}}$ models. Aggregation overhead (seconds) is measured end to end with AMC23 40 questions with k=5.
Method
Aggregation
Overhead (s)
Datasets Avg
@@
USC w/ Qwen 7B 5.89 84.15 51.6 6.67 37.5 19.29 39.84
LLama ORM (8B) 0.61 89.61 51.8 13.33 30.00 18.4 40.63
Qwen PRM (7B) 0.64 91.51 56.20 13.33 35.00 20.77 43.36
[1pt/1pt] SSA RL (0.5B) 0.20 88.17 52.80 10.00 30.00 20.18 40.23
SSA RL (1.5B) 0.33 88.48 56.60 10.00 27.50 20.47 40.61
SSA RL (3B) 0.55 89.08 57.80 10.00 32.50 20.62 42.00
Generalization to harder datasets. In Table 1, while we have trained the SSA models with problems only from GSM8K and MATH, we observe substantial performance gains on other harder datasets (AMC23 and Olympiad). Such results indicate SSA can generalize to unseen test sets outside of the training datasets.
In addition, we tested SSA on general tasks (specifically ARC-C, MMLU-Pro, and TruthfulQA (Clark et al., 2018; Wang et al., 2024c; Lin et al., 2022)) without additional training. We observe minor gains compared to majority vote, especially when using both SFT and RL, in ARC-C and MMLU-Pro dataset, but very little for TruthfulQA dataset. Further research is needed to study generalization across different domains. More details can be found in Appendix A.8.
5.3 Comparison with Sequential Test-time Scaling
We have demonstrated the advantageous of SSA over diverse parallel scaling methods in early §5.1. We now compare the performance on SSA and reasoning models that are enabled to spend more test-time compute sequentially with an RL-based approach as in (DeepSeek-AI et al., 2025). In particular, we use the results reported by Simple-RL Zoo from Zeng et al. (2025a) because they also uses Qwen-series models and train them on GSM8K and MATH dataset. In our comparison, we match the test-time computation, specifically, our 3B SSA is trained with Qwen-2.5-7B answers, but we apply it on top of 7B, 14B, 32B $\text{LLM}_{\text{ans}}$ models. We sample 8 candidate solutions from the $\text{LLM}_{\text{ans}}$ model with a maximum token length of 1,024 per answer, roughly matching the maximal generation length, 8,192, of these reasoning models.
@@ 5.3 Comparison with Sequential Test-time Scaling
5.4 Scaling SSA to handle a larger sample set
Table 3: Performance of SSA (accuracy, % averaged over 5 datasets) under increasing number of samples $K$. * denote runs that use the adapted two-stage SSA (applied when K>=32K{>=}32). SSA at $K=15,64$ even outperform majority voting and PRM at $K=128$.
Answer K 15 32 64 128
Majority Vote 54.73 54.07 54.43 55.11
Qwen PRM (7B) 56.02 55.37 58.49 56.58
SSA RL (3B) 59.12 58.42 59.78 58.68
In this section, we study SSA performance at larger $K$ values (KK=32–128). So far, SSA considers all $K$ answers as a concatenated single input. However, concatenation at larger $K$ quickly runs into context length limitations, leading to performance degradation or failures in answer extraction. For instance, with $K=32$, the concatenated input length approaches $\sim$30K tokens, nearly saturating the 32K context window of our backbone model (Qwen-2.5-3B). Furthermore, prior work highlights that the effective usable window is often shorter than the nominal size (Yang et al., 2025a; Liu et al., 2024; Hsieh et al., 2024; Ye et al., 2025).
Two-Stage SSA
To address this, we introduce a simple yet effective two-stage adaptation of SSA for large $K$ similar to  (Toshniwal et al., 2025). In the first stage, we evenly split the $K$ samples into $l_{2}$ groups of size $l_{1}$, and run SSA independently on each group to produce $l_{2}$ intermediate candidates. In the second stage, SSA is applied to aggregate these $l_{2}$ candidates into a final answer. Instead of running SSA once, with this two-stage adaptation, we run SSA for $l_{2}$ + 1 times, $l_{2}$ times to aggregate $l_{1}$ inputs, and once to aggregate $l_{2}$ inputs. We include full descriptions of two stage SSA in Appendix A.5. For our experiments, we set the hyperparameter $l_{1}=15$. This sets $l_{2}=3$ for $K=32$, and $l_{2}=5$ for $K=64$ and $l_{2}=9$ for $K=128$.
Results Table 3 presents results of SSA under increasing $K$ on Qwen-2.5-7B outputs, compared against majority voting and PRM-7B. We find that SSA makes more efficient use of samples, achieving strong performance. For example, SSA at $K=15$ already surpasses both majority vote and PRM at $K=128$. By contrast, PRM shows less stable performance (with noticeable fluctuations across $K$), as it operates by scoring individual answers without cross-sample aggregation, making it more vulnerable to the noisy reward model.
@@ 6.2 To think or not to think
A distinct pattern we noted in our RL training is the reduction of the thinking tokens. Figure 4(b) shows how response length quickly drops during training across all model sizes. The model often simply repeats the provided instruction format for thinking (e.g., ‘<think>reasoning process here</think>’) followed immediately with the final answer. Examples are shown in Appendix C Figure 14. This contrasts with other RL-based reasoning models that generate longer explanations with more training (DeepSeek-AI et al., 2025; Zeng et al., 2025a). It is likely because our SSA is conditioned on multiple candidate solutions, reducing the utility of detailed reasoning.
@@ 7 Conclusion
Limitations
Across benchmarks, SSA succeeds mainly by picking an correct candidate among input candidate answers. Failures happens mainly when the gold answer is absent among the candidates (Appendix A.6 quantifies this trend). We experimented with enabling SSA to generate new final answers by cutting the last 10% of candidate answer tokens, but this does not yield better performance (See Appendix A.7 for details).
Future work
We outline few possibilities for further improving and extending our approach. Promising directions include scaling the number of outputs to be aggregated, as well as building SSA that can incorporate outputs from multiple LLMs. Improving its performance for diverse application beyond mathematical reasoning, as well as enhacing SSA’s new target answer synthesis ability can be fruitful. Overall, we believe SSA’s hybrid approach offers a promising direction for future research in LLM reasoning.
Acknowledgments
The work is supported by the National Science Foundation (NSF) through Awards #2131186 (CISE-MSI), #1827505 (PFI), and the US Air Force Office of Scientific Research (AFOSR) via Award #FA9550-21-1-0082. The work is also supported by a College-wide Research Vision (CRV) Fund from the CCNY Provost’s Office and the Google CyberNYC Initiative. This work used Google Cloud through the CloudBank project, which is supported by NSF Award #1925001. We gratefully acknowledge use of the research computing resources of the Empire AI Consortium, Inc, with support from Empire State Development of the State of New York, the Simons Foundation, and the Secunda Family Foundation (Bloom et al., 2025).
@@ A.1 SSA Results
In addition, we report all benchmark breakdon performance of SSA compared to the PRM 7B and majority vote in Table 4.
@@
We also report the statistical significance of all results in Table 5. For statistical testing, we employed McNemar’s test, which is appropriate for paired binary outcomes. On our combined dataset analysis (aggregating all 2563 test examples across five benchmarks), our improvements over the baseline are statistically significant (p<0.05) for all configurations.
@@ A.4 Increasing $k$ during training
A.5 Scaling K during inference and Two-stage SSA
To handle large $K$ under context limits, we use a simple two-stage adaptation of SSA.
Stage 1. We form $\ell_{2}$ overlapping groups by taking evenly spaced, cyclic windows of length $\ell_{1}$ over the $K$ candidates. Concretely, with start indices $s_{i}=\big\lfloor\tfrac{iK}{\ell_{2}}\big\rfloor$ for $i=0,\dots,\ell_{2}-1$, the $i$-th group is $G_{i}=\big[\,a_{(s_{i}+t)\bmod K}\,\big]_{t=0}^{\ell_{1}-1}$. We run SSA independently on each $G_{i}$ to produce one intermediate winner.
Stage 2. We concatenate the $\ell_{2}$ intermediate winners and run SSA once more to produce the final answer.
This makes $\ell_{2}{+}1$ SSA calls: $\ell_{2}$ calls on inputs of size $\ell_{1}$, then one call on $\ell_{2}$. We use $\ell_{1}{=}15$ in all experiments; thus $\ell_{2}{=}\lceil K/\ell_{1}\rceil$ gives $\ell_{2}{=}3$ for K=32K{=}32, $\ell_{2}{=}5$ for K=64K{=}64, and $\ell_{2}{=}9$ for K=128K{=}128. By construction, groups overlap whenever $\ell_{1}>K/\ell_{2}$ (our default), yielding a replication factor $(\ell_{1}\ell_{2})/K$ that adds controlled redundancy for robustness at modest extra compute. The details are in Algorithm 1.
In addition to the performance reported in the Table 3, we also report the aggregation overhead over the AMC23 dataset. The results are in Table 8.
Table 8: Aggregation overhead per question
Answer K 32 64 128
Qwen PRM (7B) 5.25 5.41 5.80
SSA RL (3B) 9.28 7.94 7.02
Algorithm 1 Two-stage SSA
1:Question $x$; $K$ candidates $A=\{a_{0},\dots,a_{K-1}\}$; group size $\ell_{1}$; group count $\ell_{2}$
2:if $K\leq\ell_{1}$ then
3:  return $\text{SSA}(x,A)$
4:end if
[…]
6:for $i=0$ to $\ell_{2}-1$ do $\triangleright$ Stage 1: $\ell_{2}$ calls, each on $\ell_{1}$ answers
7:  $s_{i}\leftarrow\left\lfloor\dfrac{i\cdot K}{\ell_{2}}\right\rfloor$ evenly spaced starts
8:  $G_{i}\leftarrow\big[\,a_{\,((s_{i}+t)\bmod K)}\,:\;t=0,\dots,\ell_{1}-1\big]$ cyclic window (overlap-allowed)
9:  $\hat{y}_{i}\leftarrow\text{SSA}(x,G_{i})$ infer the group’s final answer
10:  $a_{G_{i}}\leftarrow\text{MapBack}(\hat{y}_{i},G_{i})$ pick the chosen candidate’s full CoT
11:  $\texttt{stage1}.\text{append}(a_{G_{i}})$
12:end for
13:return $\text{SSA}(x,\texttt{stage1})$ $\triangleright$ Stage 2: aggregate $\ell_{2}$ winners
A.6 Error Analysis
Table 9: Accuracy and error composition for the 3 B SSA (k=5k{=}5). Columns are percentages of the whole dataset. $y^{*}\in C$ means ground-truth answer $y^{*}$ included among the 5 candidates $C$.
[…]
$y^{*}\in C$
copied
$y^{*}\notin C$
synthesized
$y^{*}\notin C$
copied
$y^{*}\notin C$
synthesized
$y^{*}\in C$
copied
$y^{*}\in C$
synthesized
GSM8K 93.3 87.7 5.5 3.3 1.0 2.3 0.2
MATH 79.2 75.6 3.6 9.4 3.6 6.6 1.2
AMC23 57.5 52.5 5.0 12.5 7.5 12.5 10.0
AIME24 13.3 13.3 0.0 23.3 50.0 3.3 10.0
Olympiad 39.8 34.9 4.9 27.6 24.6 5.6 2.4
@@ A.7 Copy or Synthesize?
A.8 Generalization to other tasks
Table 11 reports accuracy on three general tasks benchmarks. On ARC-C and MMLU-Pro the SFT + RL SSA mostly beats the majority-vote baseline (Clark et al., 2018; Wang et al., 2024c). Pure RL SSAs give smaller (sometimes negative) gains, suggesting that an SFT warm-start improves out-of-domain transfer. Pure SFT SSAs exhibit varied performance: strong at 3B scale but significantly weaker at smaller scales (0.5B and 1.5B), suggesting SFT’s generalization capability strongly depends on model capacity for out domain generalization. On TruthfulQA (truthfulness/adversarial) none of the SSAs can consistently outperform majority voting, suggesting that truthfulness might not be beneficial from answer selection. Tackling truthfulness might require different goal than reasoning among answers.
Metric Qwen2.5-7B-Instruct SSA RL SSA SFT + RL SSA SFT
pass@1 Majority Vote (5) 0.5B 1.5B 3B 0.5B 1.5B 3B 0.5B 1.5B 3B
ARC-C 88.65 91.38 88.05 91.3 88.99 91.72 91.38 91.81 69.88 90.02 92.06
A.9 Efficiency Evaluation
Compute Cost (Wall-clock)
Qwen2.5-7B 1 556.12 45.5
Majority vote (k=5k{=}5) $5\timesQwen$2.5-7B 2 780.6 49.67
Qwen PRM 7B (k=5k{=}5) $5\timesQwen$2.5-7B 21.9 2 780.6 + 21.9 = 2 802.5 52.97
SSA RL 3B (k=5k{=}5, ours) $5\timesQwen$2.5-7B 25.7 2 780.6 + 25.7 = 2 806.3 56.13
SimpleRL 7B (seq. RL) 1 pass 3 273.2 58.56
Inference Cost
To measure raw compute (flops), we follow the formula from Kaplan et al. (2020). It uses roughly $C_{forward}\approx 2N$, and the backward pass is about 2 times which is $4N$. So the total is $C\approx 6N$ FLOPs per training token, with $N$ as the model parameter and $C$ as the non embedding training compute. We use $D$ as token length.
A.10 Stability Analyses
A.10.1 Effect of Candidate Ordering
A potential concern for methods that concatenate $K$ candidates is that the relative order of those candidates might influence the SSA’s decision. To test this, we evaluate the 3B SSA (RL) under three independent random permutations of the $K$ responses of 7B answer model at inference time, holding everything else fixed. As shown in Table 14, we do not observe meaningful differences when changing the order.
A.10.2 Training different random seed
Reinforcement learning can exhibit sensitivity to initialization. We therefore train SSA 3B (RL) with another seed 16 using identical data and hyperparameters, and evaluate the resulting checkpoints under the same protocol as the main results of 7B answer model. Table 14 reports per‑$K$ accuracy. We do not see significant differences.
Table 13: Ordering sensitivity. SSA 3B (RL) evaluated with the original concatenation order vs. three random permutations at inference time. Values are accuracy (%).
Metric 5 10 15 20
Original (%) 56.13 58.53 59.12 56.78
Avg. over 3 random permutations (%) 56.06 58.70 59.40 58.47
Table 14: Training‑seed sensitivity.SSA 3B (RL) trained with different random seeds. Values are accuracy (%).
Metric 5 10 15 20
Original Random Seed 42(%) 56.13 58.53 59.12 56.78
Random Seed 16 (%) 56.09 57.27 59.49 56.64
A.11 Training Data Statistics
Table 15: Correctness distribution in the raw training pool (five sampled responses per question). Counts and column percentages are shown.
Correct GSM8K MATH Combined
0/5 618 (8.3%) 3170 (26.4%) 3788 (19.5%)
1/5 218 (2.9%) 1007 (8.4%) 1225 (6.3%)
2/5 198 (2.6%) 841 (7.0%) 1039 (5.3%)
3/5 268 (3.6%) 881 (7.3%) 1149 (5.9%)
4/5 477 (6.4%) 1157 (9.6%) 1634 (8.4%)
5/5 5694 (76.2%) 4944 (41.2%) 10638 (54.6%)
Total 7473 (100.0%) 12000 (100.0%) 19473 (100.0%)
Table 15 summarizes the distribution of instance‑level correctness in the raw training pool prior to filtering. For each question we sampled K=5K{=}5 candidate solutions using top‑$k$ sampling. “m/5m/5” indicates that exactly $m$ of the five candidates match the reference answer.
Filtering and preprocessing.
We apply two light filters before RL training:
  • Validity filter. We discard instances where more than one of the five samples is null (unparsable/empty). This removes $\approx\!0.4\%$ of GSM8K and $\approx\!9.8\%$ of MATH instances.
  • Length filter. Answer sets exhibit a long‑tailed length distribution; some concatenations exceed 8k tokens due to looping or unbounded reasoning. To control VRAM and remove pathological traces, we drop instances whose concatenated prompt + answers exceed 4k tokens. This reduces the pool from ${\sim}19$k to ${\sim}17$k instances and slightly denoises the supervision.
These filters are minimal (no step‑level labeling) and aimed purely at stabilizing training; we did not tune them for accuracy.